OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge x^3 = (9x^2+y^2)C\]

OpenStudy (helder_edwin):

you have to use implicit derivatives!

OpenStudy (lgbasallote):

this should come out as \[3(3x^2 + y^2) dx - 2xydy = 0\]

OpenStudy (lgbasallote):

that's my problem

OpenStudy (helder_edwin):

that's supposed to be the answer?

OpenStudy (lgbasallote):

yeah

OpenStudy (turingtest):

C is just a constant?

OpenStudy (lgbasallote):

yes. an arbitrary constant

OpenStudy (helder_edwin):

OK. so we have \[ \Large C=\frac{x^3}{9x^2+y^2} \] then \[ \Large 0=\frac{3x^2(9x^2+y^2)-x^3(18x+2y\frac{dy}{dx})}{?} \] then \[ \Large 0=27x^4+3x^2y^2-18x^4-2x^3y\frac{dy}{dx} \]

OpenStudy (lgbasallote):

or could be it's\[x^3 = (9x^2 +y^2)C\] that's wrong...

OpenStudy (turingtest):

I get\[3x^2dx-18Cxdx-2Cydy=0\]\[3x^3dx-18Cx^2dx-2Cxydy=0\]\[3(9x^2+y^2)dx-18Cx^2dx-2Cxydy=0\]\[3((9-18C)x^2+y^2)dx-2Cxydy=0\]now what do we put for C ?

OpenStudy (asnaseer):

if you use the method outlined by @helder_edwin you will get the right answer.

OpenStudy (lgbasallote):

how o.O

OpenStudy (helder_edwin):

then \[ \Large 0=9x^4+3x^2y^2-2x^3y\frac{dy}{dx} \] then \[ \Large 3x^2(3x^2+y^2)\,dx-2x^3y\,dy \]

OpenStudy (lgbasallote):

@TuringTest i have no value for C...it was just a general solution haha

OpenStudy (turingtest):

yeah, I couldn't simplify it, though that was my initial approach as well

OpenStudy (lgbasallote):

lol wow cool @helder_edwin !!

OpenStudy (turingtest):

well we have to get rid of it somehow, so a substitution seems to be the only way

OpenStudy (lgbasallote):

seems helder figured it out already haha

OpenStudy (turingtest):

um... is there an extra x^3 or am I misreading?

OpenStudy (lgbasallote):

divide

OpenStudy (asnaseer):

\[\begin{align} C&=\frac{x^3}{9x^2+y^2}\\ \therefore0&=\frac{3x^2}{9x^2+y^2}-\frac{x^3(18x+2yy')}{(9x^2+y^2)^2} \end{align}\]which eventually simplifies to the answer...

OpenStudy (turingtest):

oh ic

OpenStudy (turingtest):

didn't notice the x^2

OpenStudy (lgbasallote):

yup...seems turning thingies into differential equations is harder than i thought haha